OpenStudy (anonymous):

Calculate the mass (in grams) of sodium in 8.5 g Na3PO4

OpenStudy (frostbite):

Hey there. So the way we solve this problem is simply by calculating the amount of substance, \(n\), [mol] of trisodium phosphate, we then notice the ratio between sodium and trisodium phosphate is 3, therefor you need to multiple the amount of substance of trisodium phosphate by 3 to obtain the amount of substance for sodium. Finally we calculate the mass of sodium by multiplying with the molar mass of sodium. Think you can do that, or shall we look at it together?

OpenStudy (anonymous):

Is it ok if you wrote it by number,?

OpenStudy (anonymous):

@frostbite And thanks

OpenStudy (frostbite):

Else let me help a little more: We want to find the amount of substance of trisodium phosphate (Na3PO4), we do this using the relation between the molar mass and the mass of a substance: \[\large n(\text{Na}_3\text{PO}_4)=\frac{ m(\text{Na}_3\text{PO}_4) }{ M(\text{Na}_3\text{PO}_4) }\] Knowing the amount of substance for trisodium phosphate, we can calculate the amount of substance for sodium, you can imagine this reaction simply things: \[\large \text{Na}_3 \text{PO}_4 \rightarrow 3 ~ \text{Na} ~ + \text{P} + ~ 4 ~ \text{O}\] The ratio from the equation thereby becomes: \[\large \frac{ n(\text{Na}_3) }{ n(\text{Na}_3\text{PO}_4) }=3 \rightarrow n(\text{Na}_3\text{PO}_4) = \frac{n(\text{Na})}{3}\] insert into previous relation we get: \[\large \frac{ n(\text{Na}) }{ 3 }=\frac{ m(\text{Na}_3\text{PO}_4) }{ M(\text{Na}_3\text{PO}_4) }\] Using the relation between the amount of substance for sodium \( n(\text{Na})\) and the mass of sodium we can write: \[\large n(\text{Na})=\frac{ m(\text{Na}) }{ M(\text{Na}) }\] Insert into previous relation we get: \[\large \frac{ m(\text{Na}) }{ M(\text{Na}) } \frac{ 1 }{3 } = \frac{ m(\text{Na}_3\text{PO}_4) }{ M(\text{Na}_3\text{PO}_4)}\] Rearrange and solve for the mass of sodium: \[\Large m(\text{Na})=\frac{ M(\text{Na}) ~ 3 ~ m(\text{Na}_3\text{PO}_4) }{ M(\text{Na}_3\text{PO}_4) }\]

OpenStudy (frostbite):

So you know what the mass of trisodium phosphate (Na3PO4) is. And you can calculate the molar mass of sodium and trisodium phosphate like this: \[\large M(\text{Na})=22.9 ~ \frac{ \text{g} }{ \text{mol} }\] And: \[\large M(\text{Na}_3 \text{PO}_4)=3 \times M(\text{Na})+M(\text{P})+4 \times M(\text{O})\] All the M(X) can be found in the periodic table and is the mass of each element :)

OpenStudy (frostbite):

You can also look up the molar masses on the wiki.

OpenStudy (frostbite):

Use the last relation and you should be good. I just hope I could make you understand so if you face a similar problem you could solve it with no issue.

OpenStudy (anonymous):

@frostbite Thanks alot 👍🏻

OpenStudy (frostbite):

No problem at all. If you like I can offer you to try calculate the number and then I check before you hand it in.

OpenStudy (frostbite):

Got to an answer? :)

OpenStudy (anonymous):

@frostbite I appreciate it ..thank you

OpenStudy (anonymous):

😊👍🏻